# Time-dependent Predictive Accuracy Metrics in the Context of Interval Censoring and Competing Risks

### v1

@@ 4.3 Comparison between biopsy frequencies
The estimated AUC and Brier scores from two approaches and different biopsy frequencies are depicted in Figure 5, and the corresponding calculated RMSE are listed in Table 2. With shorter biopsy intervals (i.e., more frequent biopsies), the IPCW estimates of the AUC tended to have lower variability, whereas the variability in the model-based approach was less impacted by the biopsy schedules. With larger biopsy intervals, i.e., fewer biopsies, the bias in the model-based AUC estimates increased while the median of the IPCW estimates was always closer to that of the reference AUC and remained more stable across different biopsy schedules. Regarding the RMSE, the model-based approach had similar (in more frequent biopsies, e.g., random schedule from 0.3 to 1 year and that from 1 to 2 years) or better performance (in less frequent biopsy schedules, e.g., the PASS schedule and the schedule from 0.3 to 4 years) in estimating the AUC, compared to the IPCW approach. Both estimates of the Brier score were closest to the reference estimates in the most frequent biopsy schedule, and the IPCW approach outperformed the model-based approach in terms of the median estimation as well as the RMSE. In other scenarios, the model-based approach had similar or better performance in estimating the Brier score than the IPCW approach. The variability of both Brier score estimates was robust to the biopsy schedules.
@@
Figure 5: The model-based and IPCW estimations of the AUC (a) and Brier score (b) from the correctly-specified model, in the context of a PASS biopsy schedule, a random schedule from 0.3 to 1 year, a random schedule from 1 to 2 years, and a random schedule from 0.3 to 4 years, compared to the corresponding ideal scenario when there is no censoring.
@@
Model-based IPCW
PASS sched 0.047 0.053
Random sched Unif(0.3,1) 0.045 0.043
Random sched Unif(1,2) 0.047 0.046
Random sched Unif(0.3,4) 0.060 0.117
@@
Model-based IPCW
PASS sched 0.079 0.086
Random sched Unif(0.3,1) 0.056 0.039
Random sched Unif(1,2) 0.068 0.066
Random sched Unif(0.3,4) 0.098 0.135
Table 2: The root mean square error (RMSE) of the model-based and the IPCW estimations of the AUC (a) and Brier score (b) from the correctly-specified model, in the context of a PASS biopsy schedule, a random schedule from 0.3 to 1 year, a random schedule from 1 to 2 years, and a random schedule from 0.3 to 4 years, to the corresponding reference metrics without censoring.
@@ 5 Discussion
One notable point when using time-dependent accuracy metrics is the choice of the interval of interest. For the IPCW approach, it is necessary to have both cases (i.e., the interval between a negative and a positive examination, must be fully contained in the interval of interest) and controls (i.e., the negative examination is after the end of the interval of interest). Therefore, the length of the interval of interest has to be longer than the minimal length of the examination interval. In the case of a fixed schedule like the Canary PASS, the timing of the biopsies should also be taken into account. Furthermore, the model is expected to have better predictive performance for later intervals of interest since more longitudinal information is used in that case. On the other hand, when the interval of interest is late in the follow-up, there will be few(er) “absolute controls”. In our first simulation study, it was noticed that the linear model was highly comparable to the correctly-specified model. This is because the predictions were based on the PSA measurements in the first year, and the simulated (overall nonlinear) PSA trajectories were relatively linear during the first year.
@@
As with all statistical methods, the two approaches presented here have strengths and limitations. The model-based approach typically has less variability. In addition, the model-based approach is capable of handling informative censoring that depends on the history of the longitudinal outcome(s), which the IPCW approach does not directly accommodate.(Rizopoulos and Taylor, 2024) However, it tends to be overly optimistic since the estimated risk probabilities from the model itself are used in its evaluation, making this approach’s performance more dependent on the correct model specification. On the other hand, the IPCW-based AUC estimates were less biased, but this approach has the disadvantage of a potentially large information loss, particularly in the setting with interval censoring where there may be only a few “absolute cases” and/or “absolute controls”. This issue is exacerbated when the examinations are infrequent and/or the interval of interest is too long.
@@
Our evaluation of the investigated methods has some limitations. The simulation study was set up to evaluate the performance of the two methods in our own real-world application. In that setting, our choices were restricted by the fact that most of the detected progression occurred around year three and that biopsies around the most cancer progression were performed in years one, two and four. The interval of interest should start earlier than year two or end later than year four to ensure the possibility of having “absolute cases”. It is expected to improve the predictive performance if the interval of interest could be chosen to contain enough “absolute cases” and “absolute controls”, and simultaneously take advantage of enough longitudinal information. However, we added the second simulation to explore how these results can be generalized to other settings in which the intervals between examinations may be different. Moreover, in this study, we assumed the sensitivity of biopsies to be perfect, meaning that the underlying cancer progression can only occur between a negative and a positive biopsy. Nevertheless in real practice, there exist many periodic examinations with low sensitivity, indicating that the underlying event(s) might have occurred even before the negative examination. This brings extra challenge to the underlying timing of interval censored event(s), and consequently leads to difficulty in defining cases and controls. Our previous study solved this problem in the model estimation phase.(Yang et al., 2024b) Similarly, the corresponding weights in the accuracy metrics could be calculated as a weighted sum of the event-specific risks in each related examination period multiplied by the misclassification probability determined by a pre-specified test sensitivity.
@@
Data availability statement
The Canary PASS data are not publicly available due to privacy restrictions. The data can be requested from the PASS Scientific Review Committee (https://canarypass.org/researchers/).
Funding statement
The research was funded by the National Institutes of Health (the NIH CISNET Prostate Award CA253910).
Conflict of interest disclosure
The authors declare no potential conflict of interests.

### v2

@@ 4.3 Comparison between biopsy frequencies
The estimated AUC and Brier scores from three approaches and different biopsy frequencies are depicted in Figure 5, and the corresponding calculated RMSE are listed in Table 2. With shorter biopsy intervals (i.e., more frequent biopsies), the IPCW estimates of the AUC tended to have lower variability, whereas the variability in the model-based approach was less impacted by the biopsy schedules. This inflation of the variability with less frequent biopsies was observed in the naive approach as well, but was less obvious than in the IPCW approach. With larger biopsy intervals, i.e., fewer biopsies, the bias in the model-based AUC estimates increased while the median of the IPCW estimates was always closer to that of the ideal AUC and remained more stable across different biopsy schedules. All three approaches had similar RMSE of the AUC estimates in the most frequent biopsy schedule. The model-based approach, on average for the four schedules, improved the RMSE of AUC estimates by 38% relative to the IPCW approach and 10% relative to the naive approach. The model-based Brier scores were much closer to those from the ideal scenario without censoring and were more robust to lower biopsy frequencies. The Brier score estimates from the IPCW and naive approach differed a lot from those from the ideal scenario, especially with larger biopsy intervals. This tendency was much more manifest in the IPCW Brier scores. Regarding the RMSE of the Brier score, the model-based approach was on average 79% better than the IPCW approach and 76% better than the naive approach. The variability of all Brier score estimates was robust to the biopsy schedules.
@@
Figure 5: The model-based, IPCW and naive (ignoring interval censoring) estimates of the AUC (a) and Brier score (b) from the correctly-specified model, in the context of a PASS biopsy schedule, a random schedule from 0.3 to 1 year, a random schedule from 1 to 2 years, and a random schedule from 0.3 to 4 years, compared to the corresponding ideal scenario when there is no censoring.
@@
Model-based IPCW Naive
Random sched Unif(0.3,1) 0.038 0.040 0.039
Random sched Unif(1,2) 0.045 0.054 0.046
PASS sched 0.036 0.056 0.045
Random sched Unif(0.3,4) 0.047 0.118 0.054
@@
Model-based IPCW Naive
Random sched Unif(0.3,1) 0.013 0.039 0.055
Random sched Unif(1,2) 0.019 0.067 0.070
PASS sched 0.015 0.087 0.083
Random sched Unif(0.3,4) 0.023 0.135 0.084
Table 2: The root mean square error (RMSE) of the model-based, IPCW and naive (interval censoring) estimates of the AUC (a) and Brier score (b) from the correctly-specified model, in the context of a PASS biopsy schedule, a random schedule from 0.3 to 1 year, a random schedule from 1 to 2 years, and a random schedule from 0.3 to 4 years, to the corresponding reference metrics without censoring.
@@ 5 Discussion
One notable point when using time-dependent accuracy metrics is the choice of the interval of interest. For the IPCW approach, it is necessary to have both cases (i.e., the interval between a negative and a positive examination, must be fully contained in the interval of interest) and controls (i.e., the negative examination is after the end of the interval of interest). Therefore, the length of the interval of interest has to be longer than the minimal length of the examination interval. In the case of a fixed schedule like the Canary PASS, the timing of the biopsies should also be taken into account. Furthermore, the model is expected to have better predictive performance for later intervals of interest since more longitudinal information is used in that case. On the other hand, when the interval of interest is late in the follow-up, there will be few(er) “absolute controls”. In our first simulation study, it was noticed that the linear model was highly comparable to the correctly-specified model. This is because the predictions were based on the PSA measurements in the first year, and the simulated (overall nonlinear) PSA trajectories were relatively linear during the first year (see Figure S1).
@@
As with all statistical methods, the two approaches presented here have strengths and limitations. The model-based approach typically has less variability and performed much better than the IPCW approach in estimating the Brier score. In addition, the model-based approach is capable of handling informative censoring that depends on the history of the longitudinal outcome(s), which the IPCW approach does not directly accommodate.(Rizopoulos and Taylor, 2024) However, it tends to be overly optimistic since the estimated risk probabilities from the model itself are used in its evaluation, making this approach’s performance more dependent on the correct model specification. On the other hand, the IPCW-based AUC estimates were less biased, but this approach has the disadvantage of a potentially large information loss, particularly in the setting with interval censoring where there may be only a few “absolute cases” and/or “absolute controls”. This issue is exacerbated when the examinations are infrequent and/or the interval of interest is too long. The IPCW approach requires an additional assumption that the timing of the examination (in our case, the biopsies) are not dependent on any covariate information or previously observed repeated measurements. This assumption is not necessary for the model-based approach. In our application, the assumption holds, as the biopsy schedule was predefined in the Canary PASS protocol. Other than the non-parametric way to calculate IPC weights, using semiparametric models including important biomarker information as covariates is an alternative to calculate the patient-specific IPC weights.(Blanche et al., 2013)
@@
Our evaluation of the investigated methods has some limitations. The simulation study was set up to evaluate the performance of the two methods in our own real-world application. In that setting, our choices were restricted by the fact that most of the detected progression occurred around year three and that biopsies around the most cancer progression were performed in years one, two and four. The interval of interest should start earlier than year two or end later than year four to ensure the possibility of having “absolute cases”. It is expected to improve the predictive performance if the interval of interest could be chosen to contain enough “absolute cases” and “absolute controls”, and simultaneously take advantage of enough longitudinal information. However, we added the second simulation to explore how these results can be generalized to other settings in which the intervals between examinations may be different. Moreover, in this study, we assumed the sensitivity of biopsies to be perfect, meaning that the underlying cancer progression can only occur between a negative and a positive biopsy. Nevertheless in real practice, there exist many periodic examinations with low sensitivity, indicating that the underlying event(s) might have occurred even before the negative examination. This brings extra challenge to the underlying timing of interval censored event(s), and consequently leads to difficulty in defining cases and controls. Our previous study solved this problem in the model estimation phase.(Yang et al., 2024) Similarly, the corresponding weights in the accuracy metrics could be calculated as a weighted sum of the event-specific risks in each related examination period multiplied by the misclassification probability determined by a pre-specified test sensitivity.
@@
Acknowledgement
The research was funded by the National Institutes of Health (the NIH CISNET Prostate Award CA253910). The authors would also like to show our gratitude to the Canary PASS team and all study participants.
Conflict of Interest
The authors have declared no conflict of interest.